OpenStudy (anonymous):

Can you please Help me? Check my answer? Choose the correct answer. Elena bought 320 tropical fish for a museum display. She bought 7 times as many triggerfish as parrotfish. How many of each type of fish did she buy? Which is a system of equations to model the problem if x represents the number of parrotfish Elena bought and y represents the number of triggerfish she bought? A. x – y = 320 y = 7x B. x + y = 7 y = 320x C. x + 7y = 320 y = 7x D. x + y = 320 y = 7x I think it is c but I am not for sure. I still thinkin c or a

OpenStudy (texaschic101):

x = parrotfish y = triggerfish x + y = 320 --- there is a total of 320 parrotfish and triggerfish y = 7x -- 7 times as many triggerfish then parrot fish check.. x + 7x = 320 8x = 320 x = 320/8 x = 40 y = 7x y = 7(40) y = 280 so there is 280 triggerfish (y) and 40 parrotfish (x)...added together = 320

OpenStudy (anonymous):

so whats the answer im confused